# HEEPidermis: a versatile SoC for BioZ recording

### v1

@@ Abstract
Index Terms:
Bioimpdance, SoC, Open-Source, Edge AI, Biomedical SystemsI Introduction
@@ II-B Analog Front-end
VCO-based ADC A pair of Voltage-Controlled Oscillator (VCO)-based ADCs were integrated [12]. They can be used one at a time or in pseudo-differential mode. As seen in Figure 4, the VCO has an oscillation range from 36- in the range 408-. The output of one of the 31-taps of the oscillator increases a non-reset 26-bit counter that can be sampled either periodically through an integrated timer or on-demand from the CPU. The counter reading is differentiated (and subtracted, in pseudo-differential mode) before being placed into the memory-mapped output register. The outputs of each inverter of the oscillator are also exposed and can be used to compute the phase difference and gain up to 6-bit resolution in the high-performance operational region of the VCO (at low oscillation frequencies this is obscured by phase noise). The counter width was fixed at 26 bits, so the result after adding these 6-bits could fit in a 32-bit word. During impedance measurement, adjustment of the injected current DC value sets the operating point of the VCO. Higher input voltages will yield a higher gain and thus SNR, but will also drastically increase power up to , mainly due to the counter update frequency, as seen in Figure 4. Power can be reduced by an order of magnitude by lowering the operational point, and SQNR can be compensated for by increasing the oversampling ratio.
@@ III Impedance measurement process
HEEPidermis offers a versatile approach that allows the reconfiguration of the system to adapt to those needs, even in real-time, to make up for long-term variations. Two exemplary uses cases are presented to illustrate the SoC’s versatility.
@@ III-B Use case 2: Impedance measurement
| Digitization (ch; bits; rate) | 2 12-bit 0.2 mHz–10 kHz | 1 20-bit 25–64 Hz | 2 24-bit 125 Hz–8 kHz | 1 10-bit I/Q demodulator |
@@
Table I summarizes the power breakdown for impedance recording (right) at a constant sensitivity of in an average, intra-patient range of , using a current, and sampling at . A finer breakdown for this case is shown in Figure 6. With the digital back-end running at power is dominated by leakage (), and by non-clock-gated digital cells.
@@ V Conclusion
In this paper, we have presented HEEPidermis, an SoC to perform tissue impedance measurements. It integrates a versatile pair of iDACs that enable a wide range of research and therapeutic procedures. It also includes two ADC channels, allowing for a trade-off between power consumption for resolution and distortion for dynamic range. Using both in conjunction enables the adjustment of the operational point of the ADC. HEEPidermis also includes two DMA channels and an event-based sub-sampler block to free the CPU to wake up only when there are enough relevant data to process. Future work will include in-silicon testing of the SoC in a plethora of tasks to showcase its versatility in the field.

### v2

@@ Abstract
I Introduction
@@ II-B Analog Front-end
VCO-based ADC A pair of Voltage-Controlled Oscillator (VCO)-based ADCs were integrated [12]. They can be used one at a time or in pseudo-differential mode. As seen in Figure 4, the VCO has an oscillation range from 36- in the range 408-. The output of one of the 31-taps of the oscillator increases a non-reset 26-bit counter that can be sampled either periodically through an integrated timer or on-demand from the CPU. The counter reading is differentiated (and subtracted, in pseudo-differential mode) before being placed into the memory-mapped output register. The outputs of each inverter of the oscillator are also exposed and can be used to compute the phase difference and gain up to 6-bit resolution in the high-performance operational region of the VCO (at low oscillation frequencies this is obscured by phase noise). The counter width was fixed at 26 bits, so the result after adding these 6-bits could fit in a 32-bit word. The large counter size enables longer sampling periods, which improve sensitivity, but does not translate into larger ENOB. During impedance measurement, adjustment of the injected current DC value sets the operating point of the VCO. Higher input voltages will yield a higher gain and thus SNR, but will also drastically increase power up to , mainly due to the counter update frequency, as seen in Figure 4. Power can be reduced by an order of magnitude by lowering the operational point, and SQNR can be compensated for by increasing the oversampling ratio.
@@ III Impedance measurement process
The joint use of iDAC, VCO-based ADC and event-based subsampling allows HEEPidermis to be reconfigured to trade-off sensitivity, range, data rate and power, even in real-time, to make up for long-term variations. Two exemplary uses cases are presented to illustrate the SoC’s versatility.
@@ III-B Use case 2: Impedance measurement
| Digitization (ch; bits; rate) | 2 16-bit∗ 0.2 mHz–10 kHz | 1 20-bit 25–64 Hz | 2 24-bit 125 Hz–8 kHz | 1 10-bit I/Q demodulator |
@@
Table I summarizes the power breakdown for impedance recording (right) at a constant sensitivity of in an average, intra-patient range of (a dynamic range of 16 bits), using a current, and sampling at . A finer breakdown for this case is shown in Figure 6. With the digital back-end running at power is dominated by leakage (), and by non-clock-gated digital cells.
@@ V Conclusion
In this paper, we have presented HEEPidermis, an SoC to perform tissue impedance measurements. It integrates a versatile pair of iDACs that enable a wide range of research and therapeutic procedures. It also includes two ADC channels, allowing for a trade-off between power consumption for resolution and distortion for dynamic range. Using both in conjunction enables the adjustment of the operational point of the ADC. HEEPidermis also includes two DMA channels and an event-based sub-sampler block to free the CPU to wake up only when there are enough relevant data to process.